OpenStudy (anonymous):

A cone shaped hat is shown: A cone is shown with slant height 14 cm and radius 6.2 cm. The unknown height is labeled with a question mark, and the line joining the tip of the cone to the base is perpendicular to the base. What is the approximate height of the hat? 2.79 centimeters, because height = Square root of the difference of 14 and 6.2 4.49 centimeters, because height = Square root of the sum of 14 and 6.2 12.55 centimeters, because height = Square root of the difference of squares of 14 and 6.2 15.31 centimeters, because height = Square root of the sum of the squares of 14 and 6.

jimthompson5910 (jim_thompson5910):

hint 

OpenStudy (anonymous):

hold on let me see what i can do.

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

kinda like the pythagorean theorem in reverse

jimthompson5910 (jim_thompson5910):

a = 6.2 b = h c = 14 a^2 + b^2 = c^2 6.2^2 + h^2 = 14^2 solve for h

OpenStudy (anonymous):

just aminte working i love the pythagorean theorem

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

the 3rd one

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks